Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (A. Polatidis at Mon Jun 8 03:53:05 METDST 1998) Minor failures
Lost part of the FORWARD pass 3 between 1951 and 2036 UT due to a problem with the SNAP file. Otherwise no problems with the data.

Medicina (C. Bortolotti at Mon Jun 8 05:04:07 METDST 1998) Success
No known problems

Torun (K. Borkowski at Mon Jun 8 09:43:52 METDST 1998) Minor failures
There were two breaks in the recording due to strong, windy storm (once with power failure): 16:38 to 19:00 and 19:35 to 20:25.

Westerbork (Tony Foley at Mon Jun 8 11:16:44 METDST 1998) Minor failures
I had to remove scans after 00:47:40UT as the sources were out of HA range at Westerbork. Some strong EMI about 14:47UT.

Noto (C.Trigilio at Mon Jun 8 18:08:07 METDST 1998) Success

Jodrell (Alastair Gunn at Tue Jun 9 03:08:41 METDST 1998) Success
Using Mk 2 telescope. RFI in VCs 5 and 6. No other known problems.

Cambridge (Alastair Gunn at Tue Jun 9 03:09:08 METDST 1998) Minor failures
FS restarted in order to load second tape. Vacuum lost, probably at 21:43, restarted for scan at 23:18 UT. No other known problems.

Calibration (Jean\-Francois Desmurs at Thu Jun 11 12:37:53 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the mai98 directory under the name gg036.antab. Rem: Jb strange Tsys values. Cm No Tsys Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 15:01:36 METDST 1998) Minor failures
On-source times short due to lots of source-switching.